Two perfectly elastic ball, weighing 6.0 lbs and 4.0 lbs , approached each other with speed of 20 ft/s and 35 ft/s, respectively. What will be their speeds after they collide?


1
Expert's answer
2021-03-09T15:26:30-0500
"v_{1f}=\frac{m_1-m_2}{m_1+m_2}v_{1i}+\frac{2m_2}{m_1+m_2}v_{2i}\\
v_{1f}=\frac{6-4}{6+4}20-\frac{8}{6+4}35=-24\frac{ft}{s}"

"v_{2f}=\frac{2m_1}{m_1+m_2}v_{1i}+\frac{m_2-m_1}{m_1+m_2}v_{2i}\\
v_{2f}=\frac{12}{6+4}20-\frac{4-6}{6+4}35=31\frac{ft}{s}"